# Cross-Document Cross-Lingual NLI via RST-Enhanced Graph Fusion and Interpretability Prediction

### v2

@@ Abstract
Natural Language Inference (NLI) is a fundamental task in natural language processing. While NLI has developed many sub-directions such as sentence-level NLI, document-level NLI and cross-lingual NLI, Cross-Document Cross-Lingual NLI (CDCL-NLI) remains largely unexplored. In this paper, we propose a novel paradigm: CDCL-NLI, which extends traditional NLI capabilities to multi-document, multilingual scenarios. To support this task, we construct a high-quality CDCL-NLI dataset including 25,410 instances and spanning 26 languages. To address the limitations of previous methods on CDCL-NLI task, we further propose an innovative method that integrates RST-enhanced graph fusion with interpretability-aware prediction. Our approach leverages RST (Rhetorical Structure Theory) within heterogeneous graph neural networks for cross-document context modeling, and employs a structure-aware semantic alignment based on lexical chains for cross-lingual understanding. For NLI interpretability, we develop an EDU (Elementary Discourse Unit)-level attribution framework that produces extractive explanations. Extensive experiments demonstrate our approach’s superior performance, achieving significant improvements over both conventional NLI models as well as large language models. Our work sheds light on the study of NLI and will bring research interest on cross-document cross-lingual context understanding, hallucination elimination and interpretability inference. Our code and datasets are available at CDCL-NLI-link for peer review.
@@
Meng-Ying Yuan1, Wen-Hao Wang2, Zi-Xuan Wang1, Yu-Jie Huang1, Kang-Li Wei1, Li Fei1, Chong Teng1, Dong-Hong Ji1, 1School of Cyber Science and Engineering, Wuhan University, 2Zhejiang University {2019302180221,foxlf823,tengchong,dhji}@whu.edu.cn
@@ 1 Introduction
  • We propose a novel method, which leverages RST-enhanced graph fusion to align semantic concepts and discourse structures, and improves interpretability by generating extractive explanations based on EDUs.
  • We conduct extensive experiments, outperforming all the baselines by at least 3.5%.
@@ 2.2 Document-level NLI
Recent approaches emphasize discourse structure and long-range dependencies. R2F Wang et al. (2022) introduces explicit reasoning extraction, and DocInfer Mathur et al. (2022) uses hierarchical encoding to model document structure, highlighting the need to capture document-level semantics.
@@
Although prior studies have advanced sentence-level and document-level NLI, challenges in cross-document and cross-lingual NLI remain largely unaddressed. Our work fills this gap by introducing the CDCL-NLI dataset and proposing an interpretable RST-enhanced graph fusion method.
Figure 2: Overview of the CDCL-NLI dataset construction process and a data example. Premise contains $D_{1}$ and $D_{2}$. Explanation is extracted from premise to enhance interpretability.
@@ 3.1 Task Formulation
[…]
between a given premise P and hypothesis H. Specifically, the premise P consists of two documents $D_{1}$ and $D_{2}$, written in different languages but discussing the same topic. The hypothesis H is a sentence-level statement. The task requires reasoning over the combined information from P with $H$ to determine their entailment relationship, involving both cross-document and cross-lingual premise integration.
3.2 Premise Data Collection
We collect our premises from GlobeSumm Ye et al. (2024), a multi-document cross-lingual summarization dataset covering 370 topics across 26 languages. In GlobeSumm, documents for each topic span diverse media outlets, publication times, and languages, providing a rich foundation for cross-document and cross-lingual inference tasks. We curated CDCL-NLI dataset by stratified randomly selecting documents for each topic to form premise pairs. To enhance cross-lingual coverage, we strategically expanded our document collection through translation.After rigorous quality filtering, our final dataset consists of high-quality inference instances covering 26 language combinations. Detailed premise establishment criteria and quality filtering standards are provided in Appendix A.1.
@@ 3.3 Hypothesis Generation and Label Specification
For each pair, we generate hypotheses across three NLI categories. Initial hypotheses are generated by GPT-4o OpenAI (2024) following specific guidelines to ensure balanced label distribution and sufficient reasoning depth. Entailment hypotheses require joint or consistent support from the premise documents. Neutral hypotheses are plausible but neither supported nor contradicted. Contradiction hypotheses explicitly conflict, focusing on cross-document inconsistencies. To reduce hallucination, GPT-4o first generates explanations before finalizing hypotheses. Detailed prompts and protocols are included in Appendix A.2.
@@ 3.4 Manual Annotation and Quality Control
Our annotation involved two phases: hypothesis verification and EDU-based explanation (Figure 2). Three graduate students independently labeled premise-hypothesis pairs, achieving strong inter-annotator agreement (Cohen’s $\kappa$: 0.71–0.82 across classes). For explanations, annotators selected minimal EDU sets supporting their decisions, with high agreement (Jaccard: 0.91; span overlap: 0.94; conclusion: 1.00). All annotations were reconciled through discussions to ensure quality (see Appendix A.3).
@@ 4.1 RST Graph Construction and Fusion
Figure 4: Our CDCL-NLI framework processes premise (${D_{1},D_{2}}$) and hypothesis through: 1) RST Graph Construction, merging ${G_{D_{1}}}$ and ${G_{D_{2}}}$ into ${G_{P}}$; 2) Graph Representation via RST-GAT layers; 3) Interpretability and Classification, extracting node-level explanations while using $\bm{h}_{G_{p}}$ and $\bm{h}_{hypo}$ for final NLI label prediction.
@@
[…]
@@
Based on the RST tree $\mathcal{T}$, we construct graphs $G_{D_{1}}$ and $G_{D_{2}}$ for each document $D_{1}$ and $D_{2}$ respectively as shown in Figure 4. For graph $G({V},{E},{R})$, we define:
  • Node Set $\mathcal{V}=\{v_{i}\mid\text{EDU}_{[s\to t]}\in\mathcal{T}\}$, where each $v_{i}$ has features: $\text{Text}_{v_{i}}$, $\phi_{v_{i}}$, and $\text{Type}_{v_{i}}$.
  • Edge Set =$\mathcal{E}=\{(v_{i},v_{j})\mid v_{i}\neq v_{j},(v_{i},v_{j},r)\in\mathcal{T}\}$, representing bidirectional edges.
@@ 4.2 Graph Representation Generation
$\mathcal{L}_{\text{triplet}}=\max(0,d(a,p)-d(a,n)+\sigma)\\ +\max(0,d(a,\text{neu})-d(a,n)+\theta),\end{aligned}$
where $d(x,y)$ is the Euclidean distance, and $a,p,\text{neu},n$ denote the premise paired with entailment, neutral, and contradiction hypotheses, respectively. Margins $\sigma$ and $\theta$ enforce the semantic order: entailment $<$ neutral $<$ contradiction.
@@ 5.1 Experiment Settings
Conventional Model Hypothesis-only 35.78 36.02 35.84 35.89 35.97 36.12
@@
Large Language Model Llama-3-8B 45.94 52.62 48.07 51.69 57.98 53.03
@@
CDCL-NLI Model Ours 71.09 70.84 68.95 72.65 72.46 70.68
@@ 6 Conclusion
This work systematically investigates CDCL-NLI, addressing key challenges in cross-document reasoning and multilingual understanding. We introduce a novel CDCL-NLI dataset spanning 26 languages and comprising 25,410 meticulously annotated instances. And we propose an RST-enhanced graph fusion mechanism with explanation prediction. Through extensive experiments and analyses, we demonstrate that our method effectively captures both structural and semantic information across documents and languages. Specifically, the RST-enhanced graph fusion mechanism and explanation prediction component not only improve model interpretability but also enhance performance, as validated by our ablation studies. Our empirical findings provide several key insights: 1) integrating rhetorical structure significantly improves the graph model’s ability to capture document-level discourse information; 2) reasoning cross-document is necessary and our method has strong cross-lingual reasoning capability; and 3) the EDU-level attribution method has a beneficial effect on classification and could generate explanations aligned with human reasoning.
@@ A.3 Data Quality Assessment
Category Description (Metric) Init. Final
Entailment (Cohen’s $\kappa$) 0.75 0.82
NLI Label Neutral (Cohen’s $\kappa$) 0.71 0.79
Contradiction (Cohen’s $\kappa$) 0.74 0.81
EDU Selection (Jaccard Sim.) 0.76 0.91
Explanation Span Coverage (Overlap Ratio) 0.81 0.94
Explanation Consistency (Align.) 0.85 1.00
@@
  • EDU Selection achieved 91% Jaccard similarity, indicating strong consensus on evidence selection
  • Span Coverage reached 94% overlap ratio, demonstrating precise identification of relevant text spans
  • Explanation Consistency achieved perfect alignment (1.00), ensuring logical coherence in reasoning
Our annotation quality assessment demonstrated strong reliability across all NLI categories. The inter-annotator agreement measured by Cohen’s $\kappa$ showed substantial initial agreement (0.71-0.75) and improved significantly after reconciliation (0.79-0.82). Specifically:
  • Entailment labels achieved the highest final agreement ($\kappa=0.82$)
  • Contradiction cases showed strong consensus ($\kappa=0.81$)
  • Neutral instances, while slightly lower, maintained robust agreement ($\kappa=0.79$)
@@ B.4 Graph Attention Formulas
[…]
@@
Baseline Base Model Optimizer LR Batch Size Max Length Epochs
Hypothesis-only XLM-R Large AdamW $3\times 10^{-6}$ 16 512 20
DocNLI XLM-R Large AdamW $3\times 10^{-6}$ 16 512 20
R2F XLM-R Large AdamW $1\times 10^{-6}$ 16 512 20
Ours XLM-R Large AdamW $1\times 10^{-5}$ 16 512(per EDU) 20
@@ C.2 LLM Fine-tuning Hyperparameters
Appendix D Additional Experiments
D.1 LLM Few-shot Prompt
As shown in Figure13, one example is provided to demonstrate how to determine the logical relationship between the premise and the hypothesis. The model is instructed to output exactly one of three labels: entailment, contradiction, or neutral. This prompt effectively guides the model to understand the task objective and output format, thereby enhancing its reasoning capability across multiple languages and documents during the few-shot validation stage.
D.2 LLM in Zero-shot Scenario
D.3 Baseline Evaluation in Single Document Scenario
Appendix E Case Study
E.1 Our Method Case
@@
[…]
@@ E.2 LLM Answer Case
Hypothesis Generation Prompt [Hypothesis Generation Prompt] We are creating a cross-document cross-lingual NLI dataset. Below are two documents under the event topic: [CATEGORY], treated as one premise in this NLI task. Based on them, generate hypotheses in three labels. You must strictly follow the instructions: 1. Hypothesis: The hypothesis should be a factual statement based on the content of the articles. It must be a simple statement and should not contain any explanation or analysis like “this contradicts” or “this agrees with” or “this is inconsistent with.” 2. Evidence: The evidence section should explain how the hypothesis relates to the articles, including any contradictions or confirmations, using specific quotes from the articles. Document Details: Document 1: Date: [DATE_1]; Article: [ARTICLE_1] Document 2: Date: [DATE_2]; Article: [ARTICLE_2] [Task 1: Entailment Generation] Generate an Entailment Hypothesis and evidence. The hypothesis is supported if evidence from both documents together or from one document alone (without contradiction in the other) logically supports it. Guidelines: Ensure each detail is verifiable by premise Include specific facts (dates, names, etc.) No speculation—strictly based on facts Evidence: Quote relevant parts from both articles and explain how they jointly support the hypothesis [Task 2: Neutral Generation] Generate a Neutral Hypothesis and evidence. One hypothesis is neutral if there is insufficient or only partial evidence in the premise to confirm or deny it, or if it contains information beyond what the premise verify. Guidelines: Reasonable speculation or expanded related aspects in a reasonable way Propose middle ground if there’s conflicting information Evidence: Show partial support from one or both articles without full confirmation Explain how the hypothesis goes beyond but stays consistent with the Document content Remember, A neutral hypothesis should not be directly confirmed by the premise (which would make it entailed), nor should it contradict the articles (which would make it conflicting). [Task 3: Conflicting Generation] Generate a Conflicting Hypothesis and evidence. One hypothesis is contradicted if either document or their combined information directly opposes it, or if the documents conflict with each other regarding the hypothesis. Guidelines: Negate or reverse key information in premise Complex and multi-faceted hypothesis with multiple contradictions Try to combine multiple points of contradiction Ensure the hypothesis appears reasonable but actually conflicts clearly Evidence: Show which document(s) the hypothesis contradicts and explain specific points If applicable, explain why this hypothesis cannot coexist with the premise content Output in JSON format: {mdframed}[backgroundcolor=gray!10, roundcorner=2pt, linewidth=0.5pt, innertopmargin=2pt, innerbottommargin=2pt, innerleftmargin=2pt, innerrightmargin=2pt]
@@
Validation Prompt You are tasked with a cross-document and cross-language Natural Language Inference (NLI) task. Your goal is to determine the relationship between the "premise" and the "hypothesis". The premise consists of two documents presented in different languages. Here is one example: Premise(Document1 in Spanish) Al menos 27 personas murieron en Perú y otras dos fueron rescatadas luego de un incendio el sábado en una mina de oro en la sureña provincia de Condesuyos, informó el Ministerio Público de ese país. Según las primeras investigaciones, la tragedia tuvo lugar tras producirse un cortocircuito a 100 metros de la entrada de la mina Yanaquihua, conocida como Esperanza I. “Se habría producido un cortocircuito que provocó un incendio en el interior del socavón, que habría puesto en riesgo la vida de los trabajadores”, informó el Gobierno regional de Arequipa. Medios locales indicaron que 27 trabajadores atrapados habían fallecido por asfixia. La noche del sábado, el Ministerio del Interior confirmó en su cuenta de Twitter el accidente. Personal policial se encuentra en el distrito de Yanaquihua para apoyar en las labores de rescate de los cuerpos de mineros que fallecieron dentro de un socavón en la provincia de Condesuyos. Imágenes difundidas en redes sociales mostraban una gran columna de humo negro proveniente de la mina, y medios locales indicaron que en el momento del cortocircuito había personal trabajando a unos 80 metros de profundidad. Se informó que el fuego se propagó de manera muy rápida por las estructuras de madera que sostienen el yacimiento, dedicado a la extracción de oro, lo que impidió que los mineros pudieran escapar.James Casquino, alcalde de Yanaquihua, dijo que el dueño de la mina fue a la comisaría de ese distrito para pedir ayuda en el rescate de las personas que se encontraban atrapadas. Premise(Document2 in Czech) 7. května (Reuters) - Úřadníci uvedli v neděli, že nehoda v malé zlaté dolině na jihu Peru odnesla život 27 pracovníků. Jedná se o jeden z nejúmrtnějších důležitých událostí v těžebním průmyslu v tomto jihoamerickém státě. Nehoda se stala v sobotu ráno v těžební společnosti Yanaquihua, která se nachází v provincii Condesuyos v departementu Arequipa. Zdá se, že došlo ke zkratu, která způsobila požár uvnitř tunelu, uvedla regionální vláda. „Informace jsou správné, potvrdila je policie v Yanaquihuě, jsou 27 obětí,“ řekl prokurátor Giovanni Matos místní televizní stanici Canal N. Regionální vláda Arequipy a ministerstvo vnitra mobilizovaly policii, zdravotníky a sanitky, aby pomohly při péči o oběti a jejich záchraně. Podle statistik peruánského ministerstva těžeb a energie je toto nejvyšší počet obětí v jediném těžebním nehodě nejméně od roku 2000. Společnost okamžitě nereagovala na žádost o komentář. Hypothesis and Label Hypothesis: "The Yanaquihua gold mine in the province of Condesuyos, Peru, experienced a fatal incident on a Saturday morning due to a short circuit that caused a fire, leading to the death of 27 workers trapped inside the tunnel, with local authorities and rescue teams being mobilized to the site." Output: entailment Now, given the following input, output exactly one word indicating the NLI label: neutral, entailment, or contradiction. Premise:[Premsise] Hypothesis: [Hypothesis] Output the NLI label (neutral/entailment/contradiction):

### v3

@@ Abstract
Natural Language Inference (NLI) is a fundamental task in natural language processing. While NLI has developed many subdirections such as sentence-level NLI, document-level NLI and cross-lingual NLI, Cross-Document Cross-Lingual NLI (CDCL-NLI) remains largely unexplored. In this paper, we propose a novel paradigm: CDCL-NLI, which extends traditional NLI capabilities to multi-document, multilingual scenarios. To support this task, we construct a high-quality CDCL-NLI dataset including 25,410 instances and spanning 26 languages. To address the limitations of previous methods on CDCL-NLI task, we further propose an innovative method that integrates RST-enhanced graph fusion with interpretability-aware prediction. Our approach leverages RST (Rhetorical Structure Theory) within heterogeneous graph neural networks for cross-document context modeling, and employs a structure-aware semantic alignment based on lexical chains for cross-lingual understanding. For NLI interpretability, we develop an EDU (Elementary Discourse Unit)-level attribution framework that produces extractive explanations. Extensive experiments demonstrate our approach’s superior performance, achieving significant improvements over both conventional NLI models as well as large language models. Our work sheds light on the study of NLI and will bring research interest on cross-document cross-lingual context understanding, hallucination elimination and interpretability inference. Our code and dataset are available at CDCL-NLI-link.
@@
Mengying Yuan1*, Wenhao Wang2*, Zixuan Wang1, Yujie Huang1, Kangli Wei1, Fei Li1, Chong Teng1, Donghong Ji1, 1Key Laboratory of Aerospace Information Security and Trusted Computing, Ministry of Education, School of Cyber Science and Engineering, Wuhan University 2Zhejiang University {yuanmengying_51,zixuanwang_nlp,huang-yj,kangliwei,lifei_csnlp,tengchong,dhji}@whu.edu.cn
\$*\$\$*\$footnotetext: Equal contribution.footnotetext: Corresponding author.
@@ 1 Introduction
  • We propose a novel method that leverages RST-enhanced graph fusion to align semantic concepts and discourse structures. The approach also enhances interpretability by generating extractive, EDU-level explanations.
  • We conduct extensive experiments demonstrating our method’s effectiveness, outperforming all baselines by at least 3.5% and establishing a new benchmark for the CDCL-NLI task.
@@ 2.2 Document-level NLI
Recent approaches emphasize discourse structure and long-range dependencies Chen et al. (2025). R2F Wang et al. (2022) introduces explicit reasoning extraction, and DocInfer Mathur et al. (2022) uses hierarchical encoding to model document structure, highlighting the need to capture document-level semantics.
@@
2.3 Graph-based Reasoning for NLI
Leveraging graph structures for semantic reasoning has emerged as a powerful paradigm. Discourse-aware graph networks model logical relationships within text for tasks like logical reasoning Hou et al. (2022); Galitsky and Ilvovsky (2025). Similarly, AMR-based graph reasoning uses Abstract Meaning Representation (AMR) to enhance question answering by providing a structured semantic representation Huang and Zhang (2025). Furthermore, prior work on graph merging and fusion has explored combining structures like AMR, RST, and CST for tasks such as multi-document summarization and inference Banarescu et al. (2018); Liao et al. (2021); Shi et al. (2024).
Although prior studies have advanced sentence-level and document-level NLI, and graph-based methods have been applied to various reasoning tasks, the challenges in cross-document and cross-lingual NLI remain largely unaddressed. Our work fills this gap by introducing the CDCL-NLI dataset and proposing a systematic integration of an interpretable RST-enhanced graph fusion method to tackle these unique complexities.
Figure 2: Overview of the CDCL-NLI dataset construction process and a data example. Premise contains $D_{1}$ and $D_{2}$. Explanation is extracted from premise to enhance interpretability. Human annotation is based on language translated into English.
@@ 3.1 Task Formulation
$\textit{Label}\in\{\textit{"Entailment"},\textit{"Neutral"},\textit{"Contradiction"}\},$
between a given premise P and hypothesis H. Specifically, the premise P consists of two documents $D_{1}$ and $D_{2}$, written in different languages but discussing the same topic.
@@
The hypothesis H is a sentence-level statement. The task requires reasoning over the combined information from P with $H$ to determine their entailment relationship, involving both cross-document and cross-lingual premise integration.
3.2 Premise Data Collection
We collect our premises from GlobeSumm Ye et al. (2024), a multi-document cross-lingual summarization dataset covering 370 topics across 26 languages. In GlobeSumm, documents for each topic span diverse media outlets, publication times, and languages, providing a rich foundation for cross-document and cross-lingual inference tasks. We curated CDCL-NLI dataset by stratified randomly selecting documents for each topic to form premise pairs. To enhance cross-lingual coverage, we strategically expanded our document collection through translation. **To address the cross-lingual aspect of the task, we used the DeepL API to translate the original English documents from GlobeSumm into 25 target languages. This translation process ensures consistent, high-quality multilingual premises.** After rigorous quality filtering, our final dataset consists of high-quality inference instances covering 26 languages. Detailed premise establishment criteria and quality filtering standards are provided in Appendix A.1.
@@ 3.3 Hypothesis Generation and Label Specification
For each pair, we generate hypotheses across three NLI categories. Initial hypotheses are generated by GPT-4o OpenAI (2024) following specific guidelines Wang et al. (2024) to ensure balanced label distribution and sufficient reasoning depth. Entailment hypotheses require joint or consistent support from the premise documents. Neutral hypotheses are plausible but neither supported nor contradicted. Contradiction hypotheses explicitly conflict, focusing on cross-document inconsistencies. To reduce hallucination, GPT-4o first generates explanations before finalizing hypotheses. Detailed prompts and protocols are included in Appendix A.2.
@@ 3.4 Manual Annotation and Quality Control
Our annotation involved two phases: hypothesis verification and EDU-based explanation (Figure 2). All human annotation was conducted on the original English versions of the premises and hypotheses. This design choice ensures that annotators did not require multilingual capabilities, and it minimizes the language gap during the critical verification process. To assess inter-annotator agreement, we randomly divided our training data into three equal parts. Each part was independently annotated by two of our three graduate students. This setup allowed us to calculate Cohen’s $\kappa$ for each of the three annotator pairs, yielding an average $\kappa$ of 0.71 across these pairs, which indicates strong agreement. For explanations, annotators selected minimal EDU sets supporting their decisions, with high agreement (Jaccard: 0.91; span overlap: 0.94; conclusion: 1.00). All annotations were reconciled through discussions to ensure quality (see Appendix A.3). The final dataset contains multilingual premise-hypothesis pairs, NLI labels, and EDU node indices for explanation, with clear metadata indicating the source of each document.
@@ 4 Our Method: RST-enhanced Graph Fusion with EDU Level Interpretability
We employ DM-RST Liu et al. (2021) parser for discourse modeling as it offers an optimal balance between structural richness and computational feasibility. Compared to the locally-focused PDTB Prasad et al. (2008), RST’s hierarchical structure effectively captures document-level organization essential for cross-document reasoning. While SDRT Asher and Lascarides (2003) is semantically richer, its $O(n^{3})$ complexity is prohibitive for large-scale tasks. Our ablation study (Table 3) empirically validates the effectiveness of our RST parser, showing that including the RST graph module can significantly improve performance despite potential parsing errors.
@@ 4.1 RST Graph Construction and Fusion
Figure 4: Our CDCL-NLI framework processes premise documents ($D_{1},D_{2}$}) and a hypothesis through a multi-stage process: 1) RST Graph Construction, where an RST parser generates initial discourse structures ($G_{D_{1}}$} and $G_{D_{2}}$}) which are then fused into a single premise graph ($G_{P}$}) using semantic edges derived from XLM-RoBERTa embeddings; 2) Graph Representation, where the fused graph is processed by RST-GAT layers; and 3) Interpretability and Classification, which extracts node-level explanations and uses the graph representations ($\bm{h}_{G_{p}}$) and hypothesis representation ($\bm{h}_{hypo}$) to predict the final NLI label.
@@
We employ DM-RST Liu et al. (2021), a top-down multilingual document-level rhetorical structure parsing framework, to extract RST information from the premise documents. As shown in Figure 5, DM-RST generates two key features for document $D$: 1) EDU boundary indices and 2) RST tree parsing outputs. By processing these features, we get $\textit{D}=\{EDU_{1},EDU_{2},...,EDU_{n}\}$ and rhetorical structure tree $\mathcal{T}$. $EDU_{i}$} represents the $i\text{-th}$ EDU’s textual content. $\mathcal{T}$ is formally defined as:
@@
Based on the RST tree $\mathcal{T}$, we construct graphs $G_{D_{1}}$ and $G_{D_{2}}$ for each document $D_{1}$ and $D_{2}$ respectively as shown in Figure 4. For graph G(V,E,R)G({V},{E},{R}), we define:
  • Node Set $\mathcal{V}=\{v_{i}\mid\text{EDU}_{[s\to t]}\in\mathcal{T}\}$, where each $v_{i}$ has features: $\text{Text}_{v_{i}}$, $\phi_{v_{i}}$, and $\text{Type}_{v_{i}}$ (e.g., nucleus or satellite).
  • Edge Set $\mathcal{E}=\{(v_{i},v_{j})\mid v_{i}\neq v_{j},(v_{i},v_{j},r)\in\mathcal{T}\}$, representing typed, bidirectional edges with rhetorical relations.
@@ 4.2 Graph Representation Generation
$\begin{aligned} \mathcal{L}_{\text{triplet}}=\max(0,d(a,p)-d(a,n)+\sigma)\\ +\max(0,d(a,\text{neu})-d(a,n)+\theta),\end{aligned}$
where $d(x,y)$ is the Euclidean distance, and $a,p,\text{neu},n$ denote the premise paired with entailment, neutral, and contradiction hypotheses, respectively. Margins $\sigma$ and $\theta$ enforce the semantic order: entailment << neutral << contradiction.
@@ 5.1 Experiment Settings
Conventional
Model
Hypothesis-only 35.78 36.02 35.84 35.89 35.97 35.91
@@
Large
Language
Model
Llama-3-8B 45.94 52.62 48.07 51.69 57.98 53.03
@@
CDCL-NLI
Model
Ours 71.09 70.84 68.95 72.65 72.46 70.68
@@ 6 Conclusion
This work systematically investigates CDCL-NLI, addressing key challenges in cross-document reasoning and multilingual understanding. We introduce a novel CDCL-NLI dataset spanning 26 languages and comprising 25,410 meticulously annotated instances. And we propose an RST-enhanced graph fusion mechanism with explanation prediction. Through extensive experiments and analyses, we demonstrate that our method effectively captures both structural and semantic information across documents and languages. Specifically, the RST-enhanced graph fusion mechanism and explanation prediction component not only improve model interpretability but also enhance performance, as validated by our ablation studies.
Acknowledgments
This work is supported by the National Key Research and Development Program of China (No. 2022YFB3103602), the National Natural Science Foundation of China (No. 62176187, No. 62202210). This work is also supported by the open project of Sichuan Provincial Key Laboratory of Philosophy and Social Science for Language Intelligence in Special Education (No. YYZN-2023-1).
@@ A.3 Data Quality Assessment
Category Description (Metric) Score
Entailment (Cohen’s $\kappa$) 0.72
NLI Label Neutral (Cohen’s $\kappa$) 0.71
Contradiction (Cohen’s $\kappa$) 0.71
EDU Selection (Jaccard Sim.) 0.76
Explanation Span Coverage (Overlap Ratio) 0.81
Explanation Consistency (Align.) 0.85
@@
  • EDU Selection achieved 76% Jaccard similarity, indicating strong consensus on evidence selection
  • Span Coverage reached 81% overlap ratio, demonstrating precise identification of relevant text spans
  • Explanation Consistency achieved 85%, ensuring logical coherence in reasoning
Our annotation quality assessment demonstrated strong reliability across all NLI categories. Our initial inter-annotator agreement score is 0.71 and annotation quality is further improved through adjudication.
@@ B.3 Justification of the Cross-Document Edge Threshold $\delta$
Baseline Base Model Optimizer LR Batch Size Max Length Epochs
Hypothesis-only XLM-R Large AdamW $3\times 10^{-6}$ 16 512 20
DocNLI XLM-R Large AdamW $3\times 10^{-6}$ 16 512 20
R2F XLM-R Large AdamW $1\times 10^{-6}$ 16 512 20
Ours XLM-R Large AdamW $1\times 10^{-5}$ 16 512(per EDU) 20
@@ B.4 Graph Attention Formulas
where $\mathbf{W}_{r,k}$ is the trainable linear transformation matrix for relation $r$ and head $k$, $a_{r,k}^{(l)}$ is the learnable attention vector for relation $r$, head $k$, and layer $l$, $\cdot\|\cdot$] denotes vector concatenation, $\psi(\cdot)$ is the ELU activation function.
@@ C.2 LLM Fine-tuning Hyperparameters
$\scriptsize1⃝1$$^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize1⃝}}$ Úřadníci uvedli v neděli, že nehoda v malé zlaté dolině na jihu Peru odnesla život 27 pracovníků. 26 potvrdila je policie v Yanaquihuě,
@@ 17 Zdá se, že došlo ke zkratu, která způsobila požár uvnitř tunelu, 33 musíme jít
Appendix D Additional Experiments
D.1 LLM Few-shot Prompt
As shown in Figure 13, one example is provided to demonstrate how to determine the logical relationship between the premise and the hypothesis. The model is instructed to output exactly one of three labels: entailment, contradiction, or neutral. This prompt effectively guides the model to understand the task objective and output format, thereby enhancing its reasoning capability across multiple languages and documents during the few-shot validation stage Chen et al. (2024).
D.2 LLM in Zero-shot Scenario
D.3 Baseline Evaluation in Single Document Scenario
Appendix E Case Study
E.1 Our Method Case
@@ E.2 LLM Answer Case
Hypothesis Generation Prompt [Hypothesis Generation Prompt] We are creating a cross-document cross-lingual NLI dataset. Below are two documents under the event topic: [CATEGORY], treated as one premise in this NLI task. Based on them, generate hypotheses in three labels. You must strictly follow the instructions: 1. Hypothesis: The hypothesis should be a factual statement based on the content of the articles. It must be a simple statement and should not contain any explanation or analysis like “this contradicts” or “this agrees with” or “this is inconsistent with.” 2. Evidence: The evidence section should explain how the hypothesis relates to the articles, including any contradictions or confirmations, using specific quotes from the articles. Document Details: Document 1: Date: [DATE_1]; Article: [ARTICLE_1] Document 2: Date: [DATE_2]; Article: [ARTICLE_2] [Task 1: Entailment Generation] Generate an Entailment Hypothesis and evidence. The hypothesis is supported if evidence from both documents together or from one document alone (without contradiction in the other) logically supports it. Guidelines: Ensure each detail is verifiable by premise Include specific facts (dates, names, etc.) No speculation—strictly based on facts Evidence: Quote relevant parts from both articles and explain how they jointly support the hypothesis [Task 2: Neutral Generation] Generate a Neutral Hypothesis and evidence. One hypothesis is neutral if there is insufficient or only partial evidence in the premise to confirm or deny it, or if it contains information beyond what the premise verify. Guidelines: Reasonable speculation or expanded related aspects in a reasonable way Propose middle ground if there’s conflicting information Evidence: Show partial support from one or both articles without full confirmation Explain how the hypothesis goes beyond but stays consistent with the Document content Remember, A neutral hypothesis should not be directly confirmed by the premise (which would make it entailed), nor should it contradict the articles (which would make it conflicting). [Task 3: Conflicting Generation] Generate a Conflicting Hypothesis and evidence. One hypothesis is contradicted if either document or their combined information directly opposes it, or if the documents conflict with each other regarding the hypothesis. Guidelines: Negate or reverse key information in premise Complex and multi-faceted hypothesis with multiple contradictions Try to combine multiple points of contradiction Ensure the hypothesis appears reasonable but actually conflicts clearly Evidence: Show which document(s) the hypothesis contradicts and explain specific points If applicable, explain why this hypothesis cannot coexist with the premise content Output in JSON format:
@@
Validation Prompt You are tasked with a cross-document and cross-language Natural Language Inference (NLI) task. Your goal is to determine the relationship between the "premise" and the "hypothesis". The premise consists of two documents presented in different languages. Here is one example: Premise(Document1 in Spanish) Al menos 27 personas murieron y dos fueron rescatadas tras un incendio en la mina de oro Yanaquihua, en Condesuyos, Perú. Las investigaciones apuntan a un cortocircuito ocurrido a unos 100 metros de la entrada, que provocó un fuego que se expandió rápidamente por las estructuras de madera del socavón, dificultando la salida de los trabajadores. Autoridades regionales señalaron que la mayoría de los mineros fallecieron por asfixia. El Ministerio del Interior confirmó el accidente la noche del sábado y la policía trabaja en el rescate de los cuerpos. Imágenes difundidas mostraron una densa columna de humo, mientras que medios locales informaron que había personal laborando a 80 metros de profundidad. El alcalde de Yanaquihua indicó que el dueño de la mina acudió a la comisaría para solicitar apoyo en el rescate. Premise(Document2 in Czech) 7. května (Reuters) - Úřadníci uvedli v neděli, že nehoda v malé zlaté dolině na jihu Peru odnesla život 27 pracovníků. Jedná se o jeden z nejúmrtnějších důležitých událostí v těžebním průmyslu v tomto jihoamerickém státě. Nehoda se stala v sobotu ráno v těžební společnosti Yanaquihua, která se nachází v provincii Condesuyos v departementu Arequipa. Zdá se, že došlo ke zkratu, která způsobila požár uvnitř tunelu, uvedla regionální vláda. „Informace jsou správné, potvrdila je policie v Yanaquihuě, jsou 27 obětí,“ řekl prokurátor Giovanni Matos místní televizní stanici Canal N. Regionální vláda Arequipy a ministerstvo vnitra mobilizovaly policii, zdravotníky a sanitky, aby pomohly při péči o oběti a jejich záchraně. Podle statistik peruánského ministerstva těžeb a energie je toto nejvyšší počet obětí v jediném těžebním nehodě nejméně od roku 2000. Společnost okamžitě nereagovala na žádost o komentář. Hypothesis and Label Hypothesis: "The Yanaquihua gold mine in the province of Condesuyos, Peru, experienced a fatal incident on a Saturday morning due to a short circuit that caused a fire, leading to the death of 27 workers trapped inside the tunnel, with local authorities and rescue teams being mobilized to the site." Output: entailment Now, given the following input, output exactly one word indicating the NLI label: neutral, entailment, or contradiction. Premise:[Premsise] Hypothesis: [Hypothesis] Output the NLI label (neutral/entailment/contradiction):